OpenStudy (anonymous):

A total of 25 students have entered a spelling contest. There are 6 medals for first through sixth place that will be awarded. In how many different ways can the medals be awarded? A. 127,512,000 B. 177,100 C. 98,200 D. 101,314,000

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 25~P~6 =25!~/~(25-6)!=25!/19!=20\times 21 \times 22 \times 23 \times 24 \times 25=?}\)